# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ Appendix D Proof of Theorem 2
This appendix finished the proof of Theorem 2 in Sec. 5.3. We utilize Theorem 2, proved in Appendix B, to derive a closed-form upper bound on the probability of a safety violation for confirmation by depth $k$. We produce two upper bounds, both of which decay exponentially with $k$. The first bound has a weaker exponent, but it is easy to verify its decay within the fault tolerance (1). The second bound has a steeper exponent whose expression is more complicated.
D.1. Upper Bounding $Y$
@@
(141) $\leq P\!\big{(}L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}$
(142) $\leq P\!\big{(}L_{s}-k+A_{s,\tau+2\Delta}+\sup_{c\geq 0}\{A_{\tau+2\Delta,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}.$
@@
(144) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,\tau+2\Delta}+M>-n\right).$
Let the first H-block mined after $s+3\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let us also define $s+T+2\Delta$ to be the first time $k$ new pacers are mined. Since the height of the $k$-th new pacer is at least $h_{b}+k$, we have $s+T+2\Delta\geq\tau$ by $\tau$’s definition. We then relax (144) to
(145) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,s+T+4\Delta}+M>-n\right).$
@@
(146) $P(Y^{(k)}>-n)$ $\leq P\left(L+A_{0,2\Delta}-k+A_{2\Delta,T+6\Delta}+M>-n\right)$
(147) $=P\left(L+M+A_{0,T+6\Delta}-k+n>0\right).$
In (147), $L$, $M$, and $A_{0,T+4\Delta}$ are mutually independent. We shall derive their MGFs shortly and use them to upper bound the probability on the right hand side of (147). We take two separate approaches: In Appendix F, we let $n=ck$ with a specific choice of $c>0$ and establish Corollary 3 by showing that as long as (1) holds, (39) and (147) both vanish as $k\to\infty$.
D.2. Chernoff Bound
@@
(148) $P(X^{(k)}\geq n)$ $\leq P(X\geq n)$
(149) $\leq(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}$
@@
(150) $P(Y^{(k)}>-n)$ $\leq P(M+L+A_{0,T+6\Delta}+n-k>0)$
(151) $\leq e^{(e^{\nu}-1)6\bar{a}}\left(q(\nu)\right)^{k}e^{\nu n}E[e^{\nu M}]E[e^{\nu L}]$
@@
(154) $\gamma(\alpha,\eta)=\frac{1}{2}\left(2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}\right).$
@@
(155) $P(X^{(k)}\geq\zeta k)\leq(1-\delta)e^{\zeta k\,\ln\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)}.$
@@
(156) $E[e^{\nu^{*}L}]$ $=E[e^{\nu^{*}M}]$
(157) $=\mathcal{E}\left(1+\frac{\gamma(\bar{a},\bar{h})}{\bar{a}}\right).$
@@
(159) $f(\alpha,\eta)=\frac{e^{6\gamma}\gamma^{2}(\alpha-\eta+\alpha\eta)^{2}}{((\alpha+\gamma)(\gamma-\eta)+\alpha\eta e^{\gamma})^{2}}$
@@
(160) $\xi(\alpha,\eta)$ $=-\gamma-\ln\frac{\eta\alpha}{(\alpha+\gamma)(\eta-\gamma)}$
(161) $=\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma.$
@@
(162) $\zeta\nu^{*}-\xi(\bar{a},\bar{h})=\zeta\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}).$
@@
(163) $\zeta=\frac{\xi(\bar{a},\bar{h})}{\ln(1+\frac{\gamma}{\bar{a}})-\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})},$
@@
(166) $c(\bar{a},\bar{h})$ $=\xi(\bar{a},\bar{h})-\nu^{*}\zeta$
(167) $=\xi(\bar{a},\bar{h})(1-\ln(1+\frac{\gamma}{\bar{a}})(\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}))^{-1})^{-1},$
@@
(168) $b(\bar{a},\bar{h})$ $=f(\bar{a},\bar{h})+1-\delta$
(169) $=\frac{e^{8\gamma}\gamma^{2}(\bar{a}-\bar{h}+\bar{a}\bar{h})^{2}}{((\bar{a}+\gamma)(\gamma-\bar{h})+\bar{a}\bar{h}e^{\gamma})^{2}}+1-\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)e^{-\bar{a}}.$
@@ Appendix E Bounding the Lead
The lead increases by at most one upon the mining of an A-block and decreases by at least one upon the mining of a jumper unless the lead is already zero, in which case it stays at 0 with the mining of a jumper. Note that $L_{0}=0$. On every sample path, there exists $u\in[0,s]$ with $L_{u}=0$ such that
(171) $L_{s}-L_{u}$ $\leq A_{u,s}-J_{u,s}.$
@@
(172) $L_{s}$ $\leq\sup_{q\in[0,s]}\{A_{q,s}-J_{q,s}\}$
(173) $\leq\sup_{q\in[0,s]}\{A_{q,s}-P_{q+\Delta,s}\}$
(174) $\leq\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{s-r,s}\}+A_{s-2\Delta,s}.$
@@
Let us introduce a reverse pacer process, $P^{s}_{0,r}$, which denotes the number of pacers mined during $(s-r,s]$, i.e.,
(175) $P^{s}_{0,r}=P_{s-r,s}.$
Due to the nature of the pacer process, one can show that $P^{s}_{0,r}$ stochastically dominates $P_{0,r-\Delta}$ as a process. Since $A$ is a memoryless Poisson process, $A_{s-\Delta-r,s-2\Delta}$ is statistically the same process as $A_{0,r-\Delta}$. Recall also $A^{\prime}$ is statistically identical to $A$. We can use (174) to show that $L_{s}$ is stochastically dominated in the sense that for every real number $x$,
@@
(176) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{0,r-\Delta}\}+A_{s-2\Delta,s}\geq x)$
(177) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{0,r-\Delta}-P_{0,r-\Delta}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
(178) $\leq P(\sup_{q\geq 0}\{A_{0,q}-P_{0,q}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
(179) $=P(L+A^{\prime}_{t,t+2\Delta}\geq x).$
@@ Appendix F Proof of Corollary 3
This proof follows Appendix D.1. It suffices to show that for every $\epsilon>0$, there exists $c>0$ such that
(180) $P(X^{(k)}\geq ck)$ $\leq\epsilon/2$
(181) $P(Y^{(k)}>-ck)$ $\leq\epsilon/2$
@@
(182) $P(Y^{(k)}>-ck)$ $\leq P(L+M+A_{0,T+6\Delta}>(1-c)k)$
(183) $\leq E[e^{\nu(M+L+A_{0,T+6\Delta}+(c-1)k)}]$
(184) $=E[e^{\nu(A_{0,T+6\Delta}+(c-1)k)}]E[e^{\nu M}]E[e^{\nu L}]$
@@
Only the first term on the right hand side of (184) depends on $k$. Given $t\geq 0$, $A_{0,t}$ is a Poisson random variable with mean $at$, whose MGF is
(185) $E[e^{\nu A_{0,t}}]=e^{(e^{\nu}-1)at}.$
@@
(186) $E[e^{\eta Z}]=\frac{1}{1-\eta/h}.$
Using (185) and (186), we can write
$E[$ $e^{\nu(A_{0,T+4\Delta}+(c-1)k)}]$
(187) $=E[E[e^{\nu(A_{0,T+6\Delta})}|T]]\cdot e^{\nu(c-1)k}$
(188) $=E[e^{(e^{\nu}-1)a(T+6\Delta)}]\cdot e^{\nu(c-1)k}$
(189) $=E[e^{(e^{\nu}-1)aT}]\cdot e^{(e^{\nu}-1)6a\Delta}\cdot e^{\nu(c-1)k}$
(190) $=\left(E[e^{(e^{\nu}-1)aZ}]\right)^{k}\cdot e^{(e^{\nu}-1)(6+k)a\Delta}\cdot e^{\nu(c-1)k}$
(191) $=e^{(e^{\nu}-1)6a\Delta}\left(\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}\right)^{k}$
(192) $=e^{(e^{\nu}-1)6a\Delta}(g(\nu,c))^{k}$
@@
(193) $g(\nu,c)$ $=\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}$
(194) $=\frac{e^{(e^{\nu}-1)\bar{a}}e^{\nu(c-1)}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}$
where $\bar{a}=a\Delta$ and $\bar{h}=h\Delta$.
In the remainder of this appendix, we will specify $\nu$ and $c$ as functions of $(\bar{a},\bar{h})$ which satisfy
(195) $g(\nu,c)<1,$
so that the right hand side of (192) and thus also of (184) vanishes as $k\to\infty$. Throughout this appendix, we denote
(196) $c=\frac{1}{4}\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)$
(197) $\theta$ $=\bar{a}(e^{\nu}-1).$
We have $c>0$ for mining rate and propagation delay satisfy the fault tolerance as (1). It is also quite obvious that $12c<$ 1. Now we choose
(198) $\theta=\frac{1}{2}\min\{1,\frac{\bar{a}}{2},\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}\}$
which is positive and yields the corresponding $\nu=\ln(1+\frac{\theta}{a})>0$. We rewrite (193) into
(199) $g(\nu,c)=\frac{e^{\theta}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
By (198), we have $\theta<1$ and $\frac{\theta}{\bar{a}}<\frac{1}{2}$, so that
(200) $g(\nu,c)$ $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
(201) $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+(1-2c)\frac{\theta}{\bar{a}})}$
(202) $=\frac{1+\theta+\theta^{2}}{1+\theta+\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}}$
where we have used the fact that $(1-2c)/\bar{a}-1/\bar{h}=1+2c/\bar{a}$. Since $\theta$ is strictly less than $\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}$, we have
(203) $\theta^{2}<\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}.$
Together, (202) and (203) guarantee (195). With the chosen $\nu$ and $c$, for sufficiently large $k$, (181) holds. The proof of Corollary 3 is thus complete.
@@ Appendix G Proof of Theorem 5
(207) $F_{2}(t-k\Delta;k,h),$
where $F_{2}$ denotes the cdf of Erlang distribution. Now we only have to calculate the last term $P(A^{\prime}_{0,t+6\Delta}\leq-n+k-(i+j))$, which equals
(208) $F_{1}(-n+k-(i+j);a(t+6\Delta))).$
@@ Appendix H Proof of Theorem 6
Consider the same attack as the private-mining attack in Definition 14 except that we let the propagation delays of all H-blocks be zero before time $s-\Delta$ for simplicity. The adversary mines privately from $s-\Delta$ and delays all new H-blocks maximally. Let the first H-block mined after $s-\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let $P^{\prime}_{r}$ denote the number of new pacers mined from $s-\Delta$ up until time $s+r-\Delta$. Evidently, $P^{\prime}_{r}$ is identically distributed as $P_{0,r+\Delta}$.
A violation occurs if there exists $d\geq s+\tau$ such that 1) the lead $L_{s-\Delta}$ plus the number of A-blocks $A_{s-\Delta,d}$ is no fewer than $P^{\prime}_{d-s+\Delta}$, the number of new pacers mined between $s-\Delta$ and $d-\Delta$ (any additional ones mined after that is not yet public); and 2) at least one pacer arrives by time $d$ to carry the target transaction, i.e., $P^{\prime}_{d-s+\Delta}\geq 1$. Hence the probability of violation is lower bounded by
(212) $P$ $\left(\cup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right),$
where $\tau$ is the time it takes for $k$ such new pacers to be mined after $s$. And the probability of (212) can be further bounded from below by
$P\left(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
(213) $\quad-P\left(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
The second term is easily determined as
(214) $P(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{\tau+\Delta}=0)$
(215) $=e^{-h(\tau+2\Delta)}.$
Since $L_{s-\Delta}$, $A_{s-\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we have the following similar in Appendix D,
$P(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(216) $=P(L_{s-\Delta}+\max_{d\geq s+\tau}\{A_{s-\Delta,d}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
(217) $=P(L_{s-\Delta}+\max_{d\geq+\tau}\{A^{\prime}_{0,d-s-\Delta}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
(218) $\geq P(L_{s-\Delta}+A^{\prime}_{0,\tau-\Delta}+\max_{d\geq\tau}\{A^{\prime}_{\tau-\Delta,d-\Delta}-P^{\prime}_{0,d-\tau}\}-P^{\prime}_{\Delta}-k\geq 0)$
(219) $=P(P^{\prime\prime}_{\Delta,2\Delta}\leq A^{\prime}_{0,\tau-\Delta}+M+L_{s-\Delta}-k),$
where (218) is due to that there are $k$ new pacers during $(s,\tau]$. $M$ in (219) is defined in (143). $P^{\prime}$ and $P$ are independent identically distributed. Thus, (219) can be written as
(220) $\begin{split}P&(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})\\ &\geq 1-P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-M-L_{s-\Delta}).\end{split}$
Now, we want to calculate $P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)$ Conditioning on $P^{\prime\prime}_{\Delta,2\Delta}$,
(221) $\begin{split}P(&A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)\\ &=P(A^{\prime}_{0,\tau-\Delta}\leq k-1-L_{s-\Delta}-M)f_{1}(0,h\Delta)\\ &\;\;+P(A_{0,\tau}\leq k-L_{s-\Delta}-M)(1-f_{1}(0,h\Delta)).\end{split}$
We have
$P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$
(222) $=$ $\sum_{i=0}^{k}\sum_{j=0}^{k-i}P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)P(M=j)P(L_{s-\Delta}=i),$
where $P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$ can be calculated as

### v5

@@ Appendix D Proof of Theorem 2
This appendix finished the proof of Theorem 2 in Sec. V-C. We utilize Theorem 11, proved in Appendix B, to derive a closed-form upper bound on the probability of a safety violation for confirmation by depth $k$. We produce two upper bounds, both of which decay exponentially with $k$. The first bound has a weaker exponent, but it is easy to verify its decay within the fault tolerance (1). The second bound has a steeper exponent whose expression is more complicated.
D-A Upper Bounding $Y$
@@
$\leq P\!\big{(}L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}$ (156)
$\leq P\!\big{(}L_{s}-k+A_{s,\tau+2\Delta}+\sup_{c\geq 0}\{A_{\tau+2\Delta,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}.$ (157)
@@
$P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,\tau+2\Delta}+M>-n\right).$ (159)
Let the first H-block mined after $s$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let us also define $s+T-\Delta$ to be the first time $k$ new pacers are mined. Since the height of the $k$-th new pacer is at least $h_{b}+k$, we have $s+T-\Delta\geq\tau$ by $\tau$’s definition. We then relax (159) to
$P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,s+T+\Delta}+M>-n\right).$ (160)
@@
$P(Y^{(k)}>-n)$ $\leq P\left(L+A_{0,2\Delta}-k+A_{2\Delta,T+3\Delta}+M>-n\right)$ (161)
$=P\left(L+M+A_{0,T+3\Delta}-k+n>0\right).$ (162)
In (162), $L$, $M$, and $A_{0,T+3\Delta}$ are mutually independent. We shall derive their MGFs shortly and use them to upper bound the probability on the right hand side of (162). We take two separate approaches: In Appendix F, we let $n=ck$ with a specific choice of $c>0$ and establish Corollary 3 by showing that as long as (1) holds, (50) and (162) both vanish as $k\to\infty$.
D-B Chernoff Bound
@@
$P(X^{(k)}\geq n)$ $\leq P(X\geq n)$ (163)
$\leq(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}$ (164)
@@
$P(Y^{(k)}>-n)$ $\leq P(M+L+A_{0,T+3\Delta}+n-k>0)$ (165)
$\leq e^{(e^{\nu}-1)3\bar{a}}\left(q(\nu)\right)^{k}e^{\nu n}E[e^{\nu M}]E[e^{\nu L}]$ (166)
@@
$\gamma(\alpha,\eta)=\frac{1}{2}\left(2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}\right).$ (169)
@@
$P(X^{(k)}\geq\zeta k)\leq(1-\delta)e^{\zeta k\,\ln\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)}.$ (170)
@@
$E[e^{\nu^{*}L}]$ $=E[e^{\nu^{*}M}]$ (171)
$=\mathcal{E}\left(1+\frac{\gamma(\bar{a},\bar{h})}{\bar{a}}\right).$ (172)
@@
$f(\alpha,\eta)=\frac{e^{5\gamma}\gamma^{2}(\alpha-\eta+\alpha\eta)^{2}}{((\alpha+\gamma)(\gamma-\eta)+\alpha\eta e^{\gamma})^{2}}$ (174)
@@
$\xi(\alpha,\eta)$ $=-\gamma-\ln\frac{\eta\alpha}{(\alpha+\gamma)(\eta-\gamma)}$ (175)
$=\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma.$ (176)
@@
$\zeta\nu^{*}-\xi(\bar{a},\bar{h})=\zeta\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}).$ (177)
@@
$\zeta=\frac{\xi(\bar{a},\bar{h})}{\ln(1+\frac{\gamma}{\bar{a}})-\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})},$ (178)
@@
$c(\bar{a},\bar{h})$ $=\xi(\bar{a},\bar{h})-\nu^{*}\zeta$ (181)
$=\xi(\bar{a},\bar{h})(1-\ln(1+\frac{\gamma}{\bar{a}})(\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}))^{-1})^{-1},$ (182)
@@
$b(\bar{a},\bar{h})$ $=f(\bar{a},\bar{h})+1-\delta$ (183)
$=\frac{e^{5\gamma}\gamma^{2}(\bar{a}-\bar{h}+\bar{a}\bar{h})^{2}}{((\bar{a}+\gamma)(\gamma-\bar{h})+\bar{a}\bar{h}e^{\gamma})^{2}}+1-\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)e^{-\bar{a}}.$ (184)
@@ Appendix E Bounding the Lead
The lead increases by at most one upon the mining of an A-block and decreases by at least one upon the publication of a jumper unless the lead is already zero, in which case it stays at 0 with the publication of a jumper. Note that $L_{0}=0$. On every sample path, there exists $u\in[0,s]$ with $L_{u}=0$ such that
$L_{s}-L_{u}$ $\leq A_{u,s}-J_{u,s-\Delta}.$ (186)
@@
$L_{s}$ $\leq\sup_{q\in[0,s]}\{A_{q,s}-J_{q,s-\Delta}\}$ (187)
$\leq\sup_{q\in[0,s]}\{A_{q,s}-P_{q+\Delta,s-\Delta}\}$ (188)
$\leq\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{s-r,s-\Delta}\}+A_{s-2\Delta,s}.$ (189)
@@
Let us introduce a reverse pacer process, $P^{s-\Delta}_{0,r}$, which denotes the number of pacers mined during $(s-r,s-\Delta]$, i.e.,
$P^{s-\Delta}_{0,r}=P_{s-r,s-\Delta}.$ (190)
Due to the nature of the pacer process, one can show that $P^{s-\Delta}_{0,r}$ stochastically dominates $P_{0,r-\Delta}$ as a process. Since $A$ is a memoryless Poisson process, $A_{s-\Delta-r,s-2\Delta}$ is statistically the same process as $A_{0,r-\Delta}$. Recall also $A^{\prime}$ is statistically identical to $A$. We can use (189) to show that $L_{s}$ is stochastically dominated in the sense that for every real number $x$,
@@
$\leq P\bigg{(}\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{0,r-\Delta}\}+A_{s-2\Delta,s}\geq x\bigg{)}$ (191)
$\leq P\bigg{(}\sup_{r\in[0,s-\Delta)}\{A_{0,r-\Delta}-P_{0,r-\Delta}\}+A^{\prime}_{t,t+2\Delta}\geq x\bigg{)}$ (192)
$\leq P\bigg{(}\sup_{q\geq 0}\{A_{0,q}-P_{0,q}\}+A^{\prime}_{t,t+2\Delta}\geq x\bigg{)}$ (193)
$=P(L+A^{\prime}_{t,t+2\Delta}\geq x).$ (194)
@@ Appendix F Proof of Corollary 3
This proof follows Appendix D-A. It suffices to show that for every $\epsilon>0$, there exists $c>0$ such that
$P(X^{(k)}\geq ck)$ $\leq\epsilon/2$ (195)
$P(Y^{(k)}>-ck)$ $\leq\epsilon/2$ (196)
@@
$P(Y^{(k)}>-ck)$ $\leq P(L+M+A_{0,T+3\Delta}>(1-c)k)$ (197)
$\leq E[e^{\nu(M+L+A_{0,T+3\Delta}+(c-1)k)}]$ (198)
$=E[e^{\nu(A_{0,T+3\Delta}+(c-1)k)}]E[e^{\nu M}]E[e^{\nu L}]$ (199)
@@
Only the first term on the right-hand side of (199) depends on $k$. Given $t\geq 0$, $A_{0,t}$ is a Poisson random variable with mean $at$, whose MGF is
$E[e^{\nu A_{0,t}}]=e^{\theta t}$ (200)
where we have defined
$\theta=\left(e^{\nu}-1\right)a$ (201)
@@
$E[e^{\eta Z}]=\frac{1}{1-\eta/h}.$ (202)
Using (200), (202), and the definition of $T$, we can write
$E[$ $e^{\nu(A_{0,T+3\Delta}+(c-1)k)}]$
$=E[E[e^{\nu(A_{0,T+3\Delta})}|T]]\cdot e^{\nu(c-1)k}$ (203)
$=E[e^{\theta(T+3\Delta)}]\cdot e^{\nu(c-1)k}$ (204)
$=\left(E[e^{\theta Z}]\right)^{k}\cdot e^{\theta(3+k)\Delta}\cdot e^{\nu(c-1)k}$ (205)
$=e^{3\theta\Delta}(g(\nu,c))^{-k}$ (206)
@@
$g(\nu,c)$ $=\left(1-\frac{\theta}{h}\right)\left(1+\frac{\theta}{a}\right)^{1-c}e^{-\theta\Delta}.$ (207)
We let
$c=\frac{a}{2}\left(\frac{1}{a}-\frac{1}{h}-\Delta\right),$ (208)
which is strictly positive when (7) holds, i.e., when the parameters are within the fault tolerance. Using Taylor series expansion, we have
$g(\nu,c)$ $=1+\left(\frac{1-c}{a}-\frac{1}{h}-\Delta\right)\theta+o(\theta)$ (209)
$=1+\frac{1}{2}\left(\frac{1}{a}-\frac{1}{h}-\Delta\right)\theta+o(\theta).$ (210)
The coefficient of the first-order term in (210) is strictly positive due to (7). Hence there exists $\theta>0$ for which $g(\nu,c)>1$, so that the right hand side of (206) and thus also the right hand side of (199) vanish as $k\to\infty$. For sufficiently large $k$, (196) holds. The proof of Corollary 3 is thus complete.
@@ Appendix G Proof of Theorem 5
$F_{2}(t-k\Delta;k,h),$ (214)
where $F_{2}$ denotes the cdf of Erlang distribution. Now we only have to calculate the last term $P(A_{0,t+3\Delta}\leq-n+k-(i+j))$, which equals
$F_{1}(-n+k-(i+j);a(t+3\Delta))).$ (215)
@@ Appendix H Proof of Theorem 6
Consider the same attack as the private-mining attack in Definition 12 except that we let the propagation delays of all H-blocks be zero before time $s$ for simplicity. The adversary mines privately from $s$ and delays all new H-blocks maximally. Let the first H-block mined after $s$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let $P^{\prime}_{r}$ denote the number of new pacers mined from $s$ up until time $s+r$. Evidently, $P^{\prime}_{r}$ is identically distributed as $P_{0,r+\Delta}$.
A violation occurs if there exists $d\geq s+\tau$ such that 1) the lead $L_{s}$ plus the number of A-blocks $A_{s,d}$ is no fewer than $P^{\prime}_{d-s-\Delta}$, the number of new pacers mined between $s$ and $d-\Delta$ (any additional ones mined after that is not yet public); and 2) at least one pacer arrives by time $d$ to be the target block, i.e., $P^{\prime}_{d-s-\Delta}\geq 1$. Hence the probability of violation is lower bounded by
$P$ $\left(\cup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\geq 1\right\}\right),$ (219)
where $\tau$ is the time it takes for $k$ such new pacers to be mined after $s$. For $k\geq 1$, $P^{\prime}_{d-s-\Delta}\geq 1$. Since $L_{s}$, $A_{s,d}$ and $P^{\prime}_{d-s-\Delta}$ are independent, we have the following similar in Appendix D,
$P(\bigcup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
$=P(L_{s}+\max_{d\geq s+\tau}\{A_{s,d}-P^{\prime}_{d-s-\Delta}\}\geq 0)$ (220)
$=P(L_{s}+\max_{d\geq\tau}\{A^{\prime}_{0,d}-P^{\prime}_{d-\Delta}\}\geq 0)$ (221)
$\geq P(L_{s}+A^{\prime}_{0,\tau}+\max_{d\geq\tau}\{A^{\prime}_{\tau,d}-P^{\prime}_{0,d-\tau}\}-k\geq 0)$ (222)
$=P(0\leq A^{\prime}_{0,\tau}+M+L_{s}-k),$ (223)
where (222) is due to that there are $k$ new pacers during $(s,\tau]$. $M$ in (223) is defined in (158). $P^{\prime}$ and $P$ are i.i.d. Thus, (223) can be written as
$\begin{split}P&(\bigcup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s+\Delta}\right\})\\ &\geq 1-P(A^{\prime}_{0,\tau}\leq k-1-M-L_{s}).\end{split}$ (224)
Now, we want to calculate $P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$ We have
$P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$
$=$ $\sum_{i=0}^{k}\sum_{j=0}^{k-i}P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)P(M=j)P(L_{s}=i),$ (225)
where $P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$ can be calculated as